# Persistent Homology for Structural Characterization in Disordered Systems

### v7

@@ I Introduction
For instance, the radial distribution function (RDF) characterizes the global structure of a system by analyzing its average density distribution, which is obtained by averaging the local density of individual particles across the system [25]. Besides, in the context of feature engineering for machine learning (ML), Atom-centered symmetry functions (ACSF) [26, 27] and smooth overlap of atomic positions (SOAP) [28] characterize the local environment of individual particles using coordinate-independent functions and Gaussian smoothing combined with spherical harmonics expansion. For global characterization, the ACSF and SOAP vectors of individual particles are typically concatenated into a single vector or subjected to straightforward transformations, such as compression into vectors of uniform length. Moreover, order parameters [29, 30] are commonly used to quantify the degree of order and symmetry. For local characterization, the Steinhardt’s bond-orientational order parameter measures local structural symmetry using spherical harmonics [31], while the ten Wolde’s approach extends this by defining bonding criteria between particles and constructing bonding networks to better distinguish ordered and disordered environments [32]. For global characterization, the bond-orientational order parameters for all particles are typically averaged to quantify the global degree of order, providing an estimate of LRO formation [31, 32]. Recently, we have noted that the introduction of the perspective of inversion symmetry and its breaking [33, 34, 35, 36] has brought new insights to the relevant field. Additionally, the static structure factor (SSF) investigates multiscale order by mapping the average density distribution of particles into the frequency domain using Fourier transformations, with the ability to capture SRO and LRO by adjusting the wave vector [37]. Medium-range order (MRO) also plays a significant role in understanding the relationship between macroscopic and macroscopic properties of materials, particularly in phase behaviors [38, 39]. However, the extraction of MRO, such as local connectivity [40] or motifs [25, 41, 42], indeed still relies on averaging or statistical processing of local information of particles.
@@
The second pathway uses a traditional physics approach without ML. We analyzed the topology of both local particle environments and the global system using our proposed new metric, the Separation Index (SI). The SI maintains a comparable level of performance to that achieved by the machine learning model, aligning with the finding that high classification accuracy can be achieved with a single variable. By adjusting the neighborhood radius, SI also functions as a mechanistic model, illustrating how long-range order and global symmetry in crystalline materials emerge from the local environment of hard particles. The calculated results of SI align with those of Global Softness, offering a unified interpretation of fluidity and structural order through both ML and traditional physics perspectives.
@@ II.1.2 Homology Groups
The kernel of $\partial_{k}$, $\ker(\partial_{k})=\{\sigma\in C_{k}\mid\partial_{k}(\sigma)=0\}$, consists of $k$-chains whose boundaries are zero—these are the cycles (closed loops) in the space. All operations in $C_{k}$ are performed modulo $p$, meaning the coefficients of $k$-chains belong to the cyclic group $\mathbb{Z}_{\mathfrak{p}}$. When $\mathfrak{p}=2$, the coefficients are reduced to binary values (0 or 1), so $k$-chains and their boundaries are reduced to binary relationships, simplifying analysis by focusing on the presence (1) or absence (0) of topological features.
@@ II.1.3 Persistent Homology
where $\sigma$ is the standard deviation of the Gaussian kernel, and $\Phi(z)$ is the cumulative distribution function (CDF) of the standard normal distribution [74]. Finally, the PI is flattened into a vector for use in ML tasks.
@@ II.2.3 Identifying Rearrangements
We set $t_{R/2}=20$ and $p_{c}=0.1$ based on the assumption that most particles in a liquid state are expected to exhibit “soft” behavior. The detailed explanation of threshold selection is provided in the Appendix A.
@@ II.3.1 Basic Concepts
Firstly, data instances are transformed into a feature matrix through feature engineering, denoted as $\mathbf{X}=g(\mathbf{x})$, where $\mathbf{x}$ is a vector representing all instances. Each instance is mapped to a corresponding feature vector, forming the feature matrix $\mathbf{X}$, where the $i$-th row represents the $i$-th instance and the $j$-th column corresponds to the $j$-th feature.
@@ II.4 Descriptive Statistics on PH (non-ML approach)
As shown in Fig. 4, the PD of a homogeneous Lennard-Jones (LJ) system, generated by Protocol 1, as defined in Sec. III.1, typically exhibits a characteristic pattern:
@@ III.3 Sample Labelling
The particle samples are labelled according to Eq. 4, which is used to characterize whether the particles will undergo rearrangement within a specific time interval. The labels for particle samples are then derived from these $p_{\text{hop},p}(t)$ values, with the label for particle $p$ in the system at time step $t$ and temperature trajectory $\text{Traj}(T)$ denoted as $y^{(T)}_{p,t}\in\{\text{soft},\text{hard}\}$.
@@ III.4 Dataset Construction
MD Protocol Protocol 1 Protocol 2
Group Group 1 Group 2 Group 3
Set Set 1 Set 3 Set 2 Set 4 Not applicable
Object Particle System Particle System Not applicable
Scale Local Global Local Global Not applicable
Status collection Soft or hard Crystal, amorphous or liquid Soft or hard Crystal, amorphous or liquid Amorphous or liquid
ML task Training Testing Not applicable
Task 1 Task 2 Task 1 Task 2 Not applicable
Fig. 10a) and 10b) Fig. 10c) and 10d)
System LJ system KA system
Dynamics Isothermal evolution Linear quenching
@@ IV Experiments and Results
3. Tracking the temporal evolution of four metrics: We analyzed their variations along trajectories and compared them with $Q_{6}$ and MSD (Sec. IV.3).
@@ IV.1.1 Task Description
The motivation is clear: in liquids, most particles tend to exhibit high mobility, whereas in solids, the opposite is true. Our goal is to establish this connection through the lens of structure-function relationships in a unified mathematical framework. As a case study, we examine the relationship between particle mobility trends and the global phase structure of the system to evaluate the effectiveness of two types of descriptors.
@@ IV.2 Descriptive Statistics on PH (non-ML approach)
Therefore, beyond providing an approximate analytical mapping for classification, SI plays a more crucial role as a bridge, enabling the analysis of the relationship between the soft-hard property of a particle and the global phase structure within a unified framework. This analysis is visualized in Fig. 5.
@@
Physically, these topological differences can be understood by considering the role of cavities (mean $\beta_{2}$) and connected components (mean $\beta_{0}$). The higher mean $\beta_{2}$ peaks in hard-particle neighborhoods reflect well-defined cavities resulting from tight, symmetric packing, indicative of solid-like stability. Besides, higher mean $\beta_{0}$ values in hard-particle neighborhoods also indicate greater local connectivity, reflecting a higher number of neighboring particles. In contrast, soft-particle neighborhoods, with more dispersed and irregular cavity structures, indicate environments where particles are less tightly bound, making these neighborhoods more liquid-like. The lower connectivity in these environments, indicated by mean $\beta_{0}$, further emphasizes their liquid-like mobility.

### v8

@@ I Introduction
For instance, the radial distribution function (RDF) characterizes the global structure of a system by analyzing its average density distribution, which is obtained by averaging the local density of individual particles across the system [25]. Besides, in the context of feature engineering for machine learning (ML), Atom-centered symmetry functions (ACSF) [26, 27] and smooth overlap of atomic positions (SOAP) [28] characterize the local environment of individual particles using coordinate-independent functions and Gaussian smoothing combined with spherical harmonics expansion. For global characterization, the ACSF and SOAP vectors of individual particles are typically concatenated into a single vector or subjected to straightforward transformations, such as compression into vectors of uniform length. Moreover, order parameters [29, 30] are commonly used to quantify the degree of order and symmetry. For local characterization, the Steinhardt’s bond-orientational order parameter measures local structural symmetry using spherical harmonics [31], while the ten Wolde’s approach extends this by defining bonding criteria between particles and constructing bonding networks to better distinguish ordered and disordered environments [32]. For global characterization, the bond-orientational order parameters for all particles are typically averaged to quantify the global degree of order, providing an estimate of LRO formation [31, 32]. Recently, we have noted that the introduction of the perspective of inversion symmetry and its breaking [33, 34, 35, 36] has brought new insights to the relevant field. Additionally, the static structure factor (SSF) investigates multiscale order by mapping the average density distribution of particles into the frequency domain using Fourier transformations, with the ability to capture SRO and LRO by adjusting the wave vector [37]. Medium-range order (MRO) also plays a significant role in understanding the relationship between microscopic and macroscopic properties of materials, particularly in phase behaviors [38, 39]. However, the extraction of MRO, such as local connectivity [40] or motifs [25, 41, 42], indeed still relies on averaging or statistical processing of local information of particles.
@@
The second pathway uses a traditional physics approach without ML. We analyzed the topology of both local particle environments and the global systems using our proposed new metric, the Separation Index (SI). The SI maintains a comparable level of performance to that achieved by the machine learning model, aligning with the finding that high accuracy of phase classification can be achieved with a single variable. By adjusting the neighborhood radius, SI also functions as a mechanistic model, illustrating how long-range order and global symmetry in crystalline materials emerge from the local environment of hard particles. The calculated results of SI align with those of Global Softness, offering a unified interpretation of fluidity and structural order through both ML and traditional physics perspectives.
@@ II.1.2 Homology Groups
The kernel of $\partial_{k}$, $\ker(\partial_{k})=\{\sigma\in C_{k}\mid\partial_{k}(\sigma)=0\}$, consists of $k$-chains whose boundaries are zero—these are the cycles (closed loops) in the space. All operations in $C_{k}$ are performed modulo $\mathfrak{p}$, meaning the coefficients of $k$-chains belong to the cyclic group $\mathbb{Z}_{\mathfrak{p}}$. When $\mathfrak{p}=2$, the coefficients are reduced to binary values (0 or 1), so $k$-chains and their boundaries are reduced to binary relationships, simplifying analysis by focusing on the presence (1) or absence (0) of topological features.
@@ II.1.3 Persistent Homology
where $\sigma$ is the standard deviation of the Gaussian kernel, $\omega(b,d)=d/\max(d)$ represents the weight and $\Phi(z)$ is the cumulative distribution function (CDF) of the standard normal distribution [74]. Finally, the PI is flattened into a vector for use in ML tasks.
@@ II.2.3 Identifying Rearrangements
We set $t_{R/2}=20$ and $p_{c}=0.1$ based on the assumption that most particles in a liquid state are expected to exhibit “soft” behavior. The detailed explanation of threshold selection is provided in the Appendix A.2.
@@ II.3.1 Basic Concepts
Firstly, data instances are transformed into a feature matrix through feature engineering, denoted as $\mathbf{X}=g(\mathbf{x})$, where $\mathbf{x}$ is a vector representing all instances and $g$ represents the process of feature engineering. Each instance is mapped to a corresponding feature vector, forming the feature matrix $\mathbf{X}$, where the $i$-th row represents the $i$-th instance and the $j$-th column corresponds to the $j$-th feature.
@@ III Dataset Generation
MD Protocol Protocol 1 Protocol 2
Group Group 1 Group 2 Group 3
Set Set 1 Set 3 Set 2 Set 4 Not applicable
Object Particle System Particle System Not applicable
Scale Local Global Local Global Not applicable
State set Soft or hard Crystal, amorphous or liquid Soft or hard Crystal, amorphous or liquid Amorphous or liquid
ML task Training Testing Not applicable
Task 1 Task 2 Task 1 Task 2 Not applicable
Fig. 10a) and 10b) Fig. 10c) and 10d)
System LJ system KA system
Dynamics Isothermal evolution Linear quenching
@@ III.3 Sample Labelling
The particle samples are labelled according to Eq. 4, which is used to characterize whether the particles will undergo rearrangement within a specific time interval. The labels for particle samples are then derived from these $p_{\text{hop},p}(t)$ values, with the label for particle $p$ in the system at time step $t$ in the trajectory $\text{Traj}(T)$ denoted as $y^{(T)}_{p,t}\in\{\text{soft},\text{hard}\}$.
@@ IV Experiments and Results
3. Tracking the temporal evolution of four metrics: We analyzed the variations of SI and global softness along trajectories and compared them with $Q_{6}$ and MSD (Sec. IV.3).
@@ IV.1.1 Task Description
The motivation is clear: in liquids, most particles tend to exhibit high mobility, whereas in solids, the opposite is true.
@@ IV.2 Descriptive Statistics on PH (non-ML approach)
Therefore, beyond providing an approximate analytical mapping for three-phase classification, SI plays a more crucial role as a bridge, enabling the analysis of the relationship between the soft-hard property of a particle and the global phase structure within a unified framework. This analysis is visualized in Fig. 5.
@@
Physically, these topological differences can be understood by considering the role of cavities (mean $\beta_{2}$) and connected components (mean $\beta_{0}$). The higher mean $\beta_{2}$ peaks in hard-particle neighborhoods reflect well-defined cavities resulting from tight, symmetric packing, indicative of solid-like stability. Besides, higher mean $\beta_{0}$ values in hard-particle neighborhoods also indicate greater local connectivity, reflecting a higher number of neighboring particles. In contrast, soft-particle neighborhoods, with more dispersed and irregular cavity structures, indicate environments where particles are less tightly bound, making these neighborhoods more liquid-like. The lower local connectivity in these environments, indicated by mean $\beta_{0}$, further emphasizes their liquid-like mobility.